OpenStudy (anonymous):

integral 3dx/sqrt(1+9x^2) I just need help with part of the algebra. .

OpenStudy (anonymous):

\[\frac{ 3\sec^2\theta }{ \sqrt(1+9x^2) }\]

OpenStudy (anonymous):

dx is = 1/3 sec^2theta

OpenStudy (anonymous):

how do they get to sec^theta/sectheta

OpenStudy (anonymous):

more specifically, where does the 1/3 go.. and how do they get the 9 out of the radical while leaving the 1 there?

OpenStudy (anonymous):

\[\int\limits \frac{ 3\sec^2\theta }{ 3\sqrt{1+9x^2} }d \theta\]

OpenStudy (anonymous):

do you just cancel the 9 as a 3 and leave the sqrt 1 because sqrt 1 is 1 ?

OpenStudy (anonymous):

i guess i am getting confused at factoring the 9 out of the square root... would you have to divide the sqrt 1 by sqrt 9 as well when you factor it? usually these problems have like 9+9x^2 in the denominator, thus why i am confused.

hartnn (hartnn):

i think you are getting confused because you kept your function in both x and theta keep your function in only theta and dtheta

OpenStudy (anonymous):

sloppy typing.. they are not mixed on my paper.

OpenStudy (anonymous):